OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −3) and perpendicular to the line y = 2x + 1?

OpenStudy (anonymous):

perpendicular lines have OPPOSITE RECIPROCAL slope a --> -1/a

OpenStudy (anonymous):

so would it be -3/4?

OpenStudy (anonymous):

y = 2x + 1 slope =2 perpendicular to that line would have slope -1/2 y= -1/2 x +b plug in your points to find b

OpenStudy (anonymous):

ohh, thanks(:

OpenStudy (amistre64):

find the slope would not entail you finding the equation of the perp line ;)